# Goldfish: Monolingual Language Models for 350 Languages

### v1

@@ Abstract
For many low-resource languages, the only available language models are large multilingual models trained on many languages simultaneously. However, using FLORES perplexity as a metric, we find that these models perform worse than bigrams for many languages (e.g. 24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B). To facilitate research that focuses on low-resource languages, we pre-train and release Goldfish, a suite of monolingual autoregressive Transformer language models up to 125M parameters for 350 languages. The Goldfish reach lower FLORES perplexities than BLOOM, XGLM, and MaLA-500 on 98 of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. However, the Goldfish significantly underperform larger multilingual models on reasoning benchmarks, suggesting that for low-resource languages, multilinguality primarily improves general reasoning abilities rather than basic text generation. We release models trained on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) of text data where available. The Goldfish models are available as baselines, fine-tuning sources, or augmentations to existing models in low-resource NLP research, and they are further useful for crosslinguistic studies requiring maximally comparable models across languages.
@@
1 Introduction
Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023). For many low-resource languages, a dedicated model optimized for that language does not exist. This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022). These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020).
To address this lack of available models, we introduce Goldfish, a suite of over 1000 monolingual language models for 350 diverse languages. [^1] The name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024). The models reach lower perplexities than XGLM (Lin et al., 2022), BLOOM 7.1B (Scao et al., 2022), and MaLA-500 (Lin et al., 2024) on 98 out of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. The Goldfish also outperform simple bigram models, which are surprisingly competitive with larger models for low-resource languages (e.g. lower perplexities than BLOOM 7.1B on 43% of its languages; §4). However, despite better perplexities, the Goldfish underperform larger multilingual models on reasoning benchmarks, suggesting that multilingual pre-training may benefit abstract reasoning capabilities over more basic grammatical text generation (§5).
Data size Model output 5MB ‘‘Goldfish are a few years of the most of the most of the most...’’ 10MB ‘‘Goldfish are a great way to the best way to the best way...’’ 100MB ‘‘Goldfish are a great way to get your fish in the wild.’’ 1GB ‘‘Goldfish are a species of fish that are found in the sea.’’
@@ 1 Introduction
Figure 1: Left: Map of the 350 languages for which Goldfish models are available, using coordinates from Glottolog (Hammarström et al., 2023). Right: Sample model outputs completing the prompt ‘‘Goldfish are’’ for the eng_latn (English) model for each dataset size, using sampling temperature zero. Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4).
Finally, to enable comparisons across languages, we release monolingual models trained on comparable dataset sizes for all languages: 5MB, 10MB, 100MB, and 1GB when available, after accounting for the fact that languages require different numbers of UTF-8 bytes to encode comparable content (Arnett et al., 2024). These Goldfish serve as baselines, allowing results in diverse languages to be situated relative to comparable models. They can also be used as source models for fine-tuning or to enhance larger multilingual models in areas where those models fall short (§4). Models and code are available at https://huggingface.co/goldfish-models.
2 Related Work
@@
Low resource language modeling often leverages multilingual pre-training, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020). Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2023). However, monolingual models have still been shown to achieve better performance than multilingual models for many languages (e.g. Martin et al., 2020; Pyysalo et al., 2021; Gutiérrez-Fandiño et al., 2021; Luukkonen et al., 2023). Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020; Chang et al., 2023).
Notably, the training datasets for massively multilingual models are often heavily skewed towards high-resource languages. For example, XGLM 4.5B is trained on over 7000$\times$ more Norwegian (71GB; 5.4M native speakers) than Quechua (0.01GB; 7.3M native speakers; Lin et al., 2022; Ethnologue, 2024). In a more extreme case, BLOOM is trained on only 0.07MB of Akan (8.1M native speakers) out of 1.61TB total (4e-6% of the pre-training dataset; Scao et al., 2022). These extremely small quantities of low-resource language data often do not leverage recent efforts to compile text data in low-resource languages (Costa-jussà et al., 2022; Imani et al., 2023; Kudugunta et al., 2023), and the data imbalances are likely to severely hinder performance in low-resource languages. Indeed, we find that these models have worse perplexities than simple bigram models for many languages (§4). Unfortunately, comparable monolingual language models across many diverse languages have yet to be studied or released.
3 Models and Datasets
@@
We introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pre-trained for 350 languages. The largest model for each language is 125M parameters. We train models on 5MB, 10MB, 100MB, and 1GB of text when available after byte premium scaling (Arnett et al., 2024). Figure 1 shows a geographic map of the 350 languages, with coordinates from Glottolog (Hammarström et al., 2023), along with sample outputs from the English model for each dataset size.
@@
Table 1: Mean FLORES log-perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size. The Goldfish languages are a strict superset of these languages.
3.1 Training Datasets
We merge the massively multilingual text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language. To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022). We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data. Full dataset details are in §A.1.
To sample pre-training datasets of the desired sizes in a language $L$, we first use the Byte Premium Tool (Arnett et al., 2024) to estimate the byte premium for $L$, the number of UTF-8 bytes required to encode comparable text in $L$ relative to eng_latn (English). For example, khm_khmr (Khmer) has byte premium 3.91, meaning that it uses approximately 3.91$\times$ as many UTF-8 bytes as English to encode content-matched text. We divide each dataset size by the estimated byte premium for the corresponding language, thus measuring all datasets in units of “equivalent” English text bytes. We sample datasets to train monolingual language models on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) when available after byte premium scaling. [^2] The languages with 5MB-dataset models are a subset of the languages with 10MB-dataset models, and similarly for the 100MB and 1GB dataset sizes. These are equivalent to roughly 1M, 2M, 20M, and 200M tokens of English text respectively; including 10 epochs of repetition, the 1GB-dataset models are trained on the equivalent of roughly 2B English tokens. When a 1GB dataset is not available for a language after byte premium scaling, we include a full model (267 languages) trained on the entire dataset in that language, for use cases that seek to maximize performance in a specific low-resource language.
3.2 Architectures and Pre-Training
For each language and each dataset size, we pre-train an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019). For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019). Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes. Full hyperparameters are reported in §A.2.
We tokenize each dataset using using a monolingual SentencePiece tokenizer (Kudo and Richardson, 2018) trained on that dataset size, limiting tokenizer training text to 100MB after byte premium scaling. Following Liu et al. (2019), we use vocabulary size 50K and a maximum sequence length of 512 tokens for all models. We train each language model on 10 epochs of its corresponding dataset. [^3] Multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but we find that more than 10 epochs of training leads to overfitting for extremely small datasets (e.g. 5MB). Pre-training details, compute costs, and all available models are reported in §A.2.
@@ 3.2 Architectures and Pre-Training
# Langs Chance Goldfish XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
Belebele 121 25.0 28.2 30.1 30.6 30.2 30.6
XCOPA 11 50.0 54.9 57.9 60.6 56.9 55.6
XStoryCloze 10 50.0 52.5 57.1 59.9 58.2 55.7
Table 3: Reasoning benchmark accuracies averaged over non-English languages. Despite better perplexities, the Goldfish perform significantly worse than larger multilingual models on reasoning.
4 FLORES Log-Perplexity Evaluations
We first evaluate our models on FLORES-200 log-perplexity (Costa-jussà et al., 2022) (equivalently, negative log-likelihood; Lin et al., 2024). To avoid tokenization confounds from computing log-perplexity per token, we compute log-perplexity per FLORES sequence. Regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES. In most cases, $s$ is a single sentence. For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$. We then compute the mean over sequences:
@@ 4 FLORES Log-Perplexity Evaluations
We compare the Goldfish models with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages). We also compare to simple bigram models trained on the Goldfish datasets. [^4] Bigram and perplexity implementation details in §A.3. In all cases, we use the Goldfish model trained on the maximum amount of data in each language (maximum 1GB).
FLORES log-perplexity results.
The Goldfish reach lower log-perplexities than all four comparison models on 98 of the 204 FLORES languages. Average log-perplexities for the 100 FLORES languages included in both XGLM 4.5B and MaLA-500 are reported in Table 1 (excluding XGLM 7.5B and BLOOM 7.1B there because they are trained on far fewer languages). On average, the Goldfish reach 13% lower log-perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B.
@@
To ensure that these results are not driven by a small subset of specific languages, in Table 2 we report the pairwise “win” rates for Goldfish and bigrams vs. all four comparison models, for the set of FLORES languages shared between each pair. The Goldfish models have a perplexity win rate above 50% against all comparison models except XGLM 7.5B, which considers only 30 fairly high-resource languages (Lin et al., 2022). Notably, the bigram models also reach lower perplexities than large multilingual models for a nontrivial number of languages: 24% of languages in XGLM 4.5B and 43% of languages in BLOOM 7.1B. Still, the bigrams have worse perplexities than Goldfish for all languages. Log-perplexities for individual languages and models are reported in Table LABEL:tab:ppl-table.
5 Multilingual Reasoning Benchmarks
Because FLORES perplexities are not necessarily reflective of complex capabilities in language models, we also evaluate Goldfish, XGLM, BLOOM, and MaLA-500 (as in §4) on non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense; Lin et al., 2022). All models are evaluated zero shot with no fine-tuning. Evaluation task details are in §A.4.
Results for all three reasoning tasks are reported in Table 3. Although all models perform quite poorly (close to chance accuracy), the Goldfish perform substantially worse than the multilingual models. [^5] It is unlikely that this effect is due to model size alone; the Goldfish models (125M parameters) have easily enough capacity for their maximum of 1GB of text data. This indicates that the combination of larger datasets and model sizes in multilingual pre-training can allow language models to develop reasoning capabilities in specific languages, even when perplexities in those languages remain high. For example, XGLM 7.5B has worse perplexities than Goldfish for 82 Belebele languages (in fact, worse than bigrams for 77 languages), but it outperforms Goldfish on Belebele (reading comprehension) for 56 of those languages. This is in stark contrast with monolingual language models, which generally must reach low perplexities and acquire basic grammatical capabilities before developing reasoning abilities (Liu et al., 2021; Choshen et al., 2022; Xia et al., 2023; Chang et al., 2024). Intuitively, it may be that abstract reasoning patterns are often more language-agnostic than grammatical text generation, and thus multilingual pre-training primarily benefits the former.
6 Conclusion
We pre-train and release Goldfish, a suite of over 1000 monolingual language models for 350 languages. The Goldfish achieve perplexities that are competitive with, and on average lower than, state-of-the-art multilingual language models across languages. However, they underperform large multilingual models on reasoning tasks; in low-resource languages, it appears that multilingual pre-training facilitates nontrivial reasoning capabilities despite extremely poor perplexities. We publicly release all Goldfish models to be used as comparable baselines, fine-tuning sources, or augmentations to larger models (e.g. cross-lingual experts; Blevins et al., 2024) in future low-resource NLP research.

### v2

@@ Abstract
For many low-resource languages, the only available language models are large multilingual models trained on many languages simultaneously. Despite state-of-the-art performance on reasoning tasks, we find that these models still struggle with basic grammatical text generation in many languages. First, large multilingual models perform worse than bigrams for many languages (e.g. 24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B) using FLORES perplexity as an evaluation metric. Second, when we train small monolingual models with only 125M parameters on 1GB or less data for 350 languages, these small models outperform large multilingual models both in perplexity and on a massively multilingual grammaticality benchmark. To facilitate future work on low-resource language modeling, we release Goldfish, a suite of over 1,000 small monolingual language models trained comparably for 350 languages. These models represent the first publicly-available monolingual language models for 215 of the languages included.
Keywords: multilingual NLP, low-resource languages
[Uncaptioned image]  Models and training data       Training and evaluation code
[…]
@@
Tyler A. Chang1, Catherine Arnett2, Zhuowen Tu1,3, Benjamin K. Bergen1
1Department of Cognitive Science, 2Department of Linguistics,
3Department of Computer Science,
University of California San Diego
Abstract content
1.  Introduction
Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020b; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023). For many low-resource languages, a dedicated model optimized for that language does not exist. For example, in an analysis of publicly available models on Hugging Face, we find that 215 of the 350 languages in this paper did not have a monolingual text-generation model prior to the Goldfish models, and 47 of the languages did not have any text-generation models at all (Arnett and Chang, 2025).
This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022). These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020a).
In particular, we find in this paper that large multilingual language models still struggle with basic next token prediction—a prerequisite for text generation—in low-resource languages. To establish a baseline for basic text generation performance in low-resource languages, we introduce Goldfish [^1] This name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024). If an acronym is desired, Goldfish can stand for Generative autOregressive Low-resource moDels For lImited-compute System Hardware., a suite of over 1000 monolingual language models for 350 diverse languages.
@@ 1.  Introduction
Thus, the primary contributions of this short paper are (1) to show that large multilingual models still suffer from poor next token prediction performance in low-resource languages, (2) to demonstrate that small monolingual models often exhibit lower perplexities and more grammatical next token predictions for such languages, and (3) to address the lack of available models in low-resource languages by releasing comparable monolingual models for 350 languages. Models, training data, and code are available at: https://huggingface.co/goldfish-models.
Data size Model output
5MB Goldfish are a few years of the most of the most of the most...
10MB Goldfish are a great way to the best way to the best way...
100MB Goldfish are a great way to get your fish in the wild.
1GB Goldfish are a species of fish that are found in the sea.
@@
Figure 1: Left: Map of the 350 languages for which Goldfish models are available, using coordinates from Glottolog (Hammarström et al., 2023). Color indicates the largest dataset size for that language. Right: Sample model outputs completing the prompt ‘‘Goldfish are’’ for the eng_latn (English) model for each dataset size, using sampling temperature zero. Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4.1).
2.  Related Work
@@
Low resource language modeling often leverages multilingual pretraining, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020b). Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2024a). However, monolingual models have still been shown to achieve better performance than multilingual models for many languages (e.g. Martin et al., 2020; Pyysalo et al., 2021; Gutiérrez-Fandiño et al., 2021; Luukkonen et al., 2023). Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020b; Chang et al., 2024a).
[…]
3.  Models and Datasets
@@
To compare to larger multilingual models, we introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pretrained for 350 languages. The largest model for each language is 125M parameters. We train models on 5MB, 10MB, 100MB, and 1GB of text when available after byte premium scaling (Arnett et al., 2024). Figure 1 shows a geographic map of the 350 languages, with coordinates from Glottolog (Hammarström et al., 2023), along with sample outputs from the English model for each dataset size.
3.1.  Training Datasets
We merge the massively multilingual text datasets compiled in Chang et al. (2024a), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language. We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data. Full dataset details are in §A.1. To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022). We conduct a contamination analysis to determine whether FLORES is contaminated in our datasets and find that for 98% of languages, less than 10 out of 2000 FLORES sequences appear in the training dataset at all (§A.2).
[…]
As we cap our dataset sizes at 1GB per language, our models are trained on much less data than is otherwise available for high-resource languages. Still, when we compare our dataset sizes to FineWeb2 (Penedo et al., 2025) for the languages covered by both Goldfish and FineWeb2, 106 of those languages have more data in our dataset than in FineWeb2. We attribute this partially to our approach to data collection, which involves seeking out language-specific resources rather than relying solely on language identification models, which can be unreliable for web data in low-resource languages (Ortiz Suarez et al., 2026). We release our training datasets at: https://huggingface.co/datasets/goldfish-models/fish-food.
@@ 3.1.  Training Datasets
Table 1: Mean FLORES perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size. The Goldfish languages are a strict superset of these languages.
@@
3.2.  Architectures and Pretraining
We train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full. The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language. In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total).
For each language and each dataset size, we pretrain an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019). For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019). Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2024a), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes. Full hyperparameters are reported in §A.3.
Each model has a custom monolingual tokenizer, which is trained with a vocabulary size of 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling). We use Unigram tokenizers trained with the SentencePiece implementation (Kudo and Richardson, 2018). Training text is randomly sampled from the dataset for the desired language. [^3] To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling. After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens. We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form. Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling. [^4] When de-tokenized, the tokenized datasets are slightly smaller than the original text datasets, because the tokenizer truncates lines to create 512-token sequences. Reported dataset sizes account for truncation.
We train each model for 10 epochs of the training data; multiple epochs of pretraining is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pretraining on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios. All language model pretraining runs together take a total of $1.65\times 10^{20}$ FLOPs. This is less than $1/1900\times$ the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020a; $3.14\times 10^{23}$ FLOPs). Further pretraining details are reported in §A.3.
$3.$  Training Bigram Baselines
Because there do not exist other models for many of these languages, we also train bigram baseline models on the same training data and with the same tokenizer for each language. Each bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset. When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e. “stupid backoff”; Brants et al., 2007). We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.
4.  Evaluations
4.1.  FLORES Log-Perplexity
We first evaluate our models on FLORES-200 log-perplexity (Costa-jussà et al., 2022) (equivalently, negative log-likelihood; Lin et al., 2024). To avoid confounds from different tokenizers across models, we compute log-perplexities at the sequence level. Specifically, regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES. In most cases, $s$ is a single sentence. For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$. We then compute the mean over sequences:
@@ 4.1.  FLORES Log-Perplexity
We compare the Goldfish with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages). We also compare to our bigram baselines, which were trained on the same datasets as the Goldfish models. In all cases, we use the Goldfish model trained on the maximum amount of data in a language (maximum 1GB).
@@
The Goldfish reach lower log-perplexities than all four comparison models on 98 of the 204 FLORES languages. On average, the Goldfish reach 13% lower perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B (Table 1). To ensure that these results are not driven by a small subset of specific languages, in Table 2 we also report the pairwise “win” rates for Goldfish and bigrams vs. all four comparison models, for the set of FLORES languages shared between each pair. The Goldfish models have a perplexity win rate above 50% against all comparison models except XGLM 7.5B, which considers only 30 fairly high-resource languages (Lin et al., 2022). Notably, the bigram models also reach lower perplexities than large multilingual models for a nontrivial number of languages: 24% of languages in XGLM 4.5B and 43% of languages in BLOOM 7.1B. Still, the bigrams have worse perplexities than Goldfish for all languages. Perplexities for individual languages and models are available in our GitHub.
Dataset / Model Chance Goldfish Gemma 3 BLOOM XGLM Gemma 3 LLaMA 3 XGLM
124M 270M 560M 564M 1B 1B 1.7B
MultiBLiMP (avg) 50.0 78.8 72.3 64.2 66.6 78.0 77.4 62.7
Belebele (avg) 25.0 28.2 22.7 28.9 29.0 22.7 28.9 28.2
XCOPA (avg) 50.0 55.1 54.6 54.4 55.2 59.7 55.9 57.6
XStoryCloze (avg) 50.0 52.3 52.7 52.6 53.1 58.5 55.1 56.2
Table 3: Average performance on downstream benchmarks for all models tested.
4.2.  Downstream Tasks
Next, we evaluate linguistic knowledge (grammaticality) with MultiBLiMP (Jumelet et al., 2025) as implemented in the LM Evaluation Harness (Gao et al., 2023); this covers 74 of the Goldfish languages. We compare against popular small multilingual models: BLOOM 560M (Scao et al., 2022), XGLM (564M and 1.7B; Lin et al., 2021), Gemma 3 (270M and 1B base models; Gemma Team, 2025), and Llama 3.2 (1B base model; Meta AI, 2024). Results are in Table 3. The Goldfish models have higher average accuracy than any of the multilingual models, and they have the highest accuracy of all models tested for 25 of the 74 languages. This result highlights the benefits of small monolingual models, especially for languages which only account for a small portion of the training data in multilingual models.
We also evaluate the same models on three popular multilingual reasoning benchmarks: Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense reasoning; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense reasoning; Lin et al., 2022). All models are evaluated zero shot with log-probability solution ranking, with no fine-tuning or instruction tuning. Unfortunately, all models perform quite poorly (close to chance accuracy; Table 3), suggesting that available multilingual reasoning evaluations are not appropriate for pretrained-only models of this scale.
5.  Discussion
Our results demonstrate that large multilingual language models still struggle with basic grammatical text generation for many languages (e.g. often worse perplexities than bigrams). The Goldfish (small monolingual models) exhibit lower perplexities and more grammatical next token predictions for many low-resource languages. These results make the Goldfish suitable baselines for basic grammatical text generation in diverse languages, motivating future work developing low-resource language models. Furthermore, the Goldfish are accessible to labs with limited compute budgets, they are trained on a roughly human-like amount of data (Warstadt et al., 2023), and they are trained to be maximally comparable across languages. Future work may investigate precisely when larger-scale multilingual pretraining provides benefits to lower-resource languages; for example, it may be that abstract reasoning patterns and heuristics are often more language-agnostic than grammatical text generation, and thus larger-scale multilingual pretraining primarily benefits the former.
6.  Conclusion
In this paper, we pretrain and release Goldfish, a suite of over 1000 monolingual language models for 350 languages. For the majority of these languages, the Goldfish represent the first monolingual model dedicated for that language. The Goldfish achieve perplexities that are competitive with, and on average lower than, state-of-the-art multilingual language models across languages. However, similar to multilingual models of the same scale, the Goldfish still struggle with reasoning tasks. This suggests that smaller monolingual models may better represent linguistic knowledge of the target language, but they do not perform well on more complex tasks at this scale. We publicly release the Goldfish models to promote future research pushing the limits of language model capabilities in low-resource languages.